RST manufactures two products. Information about the two products are as follows:

Product A

Product B

Selling price per unit

$100

$50

Variable costs per unit

  $60

  $40

Contribution margin per unit

$40

$10

The company expects fixed costs to be $420,000. The firm expects 60% of its sales (in units) to be Product A (a sales mix of 3:2).

Required:

A.

Calculate the contribution margin per package.

B.

Determine the break-even point in units for Products A and B.

C.

Determine the level of sales (in dollars) necessary to generate operating income of $200,000.
Answer #1

a) Contribution margin per package = (40*60%+10*40%) = 28 per package

b) Break even point = Fixed cost/Weighted average contribution = 420000/28 = 15000 Units

Product A = 15000*60 % = 9000 Units

Product B = 15000*40% = 6000 Units

c) Weighted average ratio = (40%*60%+20%*40%) = 32%

Required sales = (420000+200000)/.32 = $1937500

Product A = 1937500*60% = 1162500

Product B = 1937500*40% = 775000